OpenStudy (anonymous):

86 is 45% of what number? (1 point) 86 = 0.45n n = 0.45 × 86 0.45 = n × 86 86 = 45n

OpenStudy (anonymous):

mutiple choice

OpenStudy (anonymous):

the first one, 86 = 0.45n divide each side by 0.45 and then it equal n

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

happy to help (:

OpenStudy (anonymous):

:) What number is 73% of 215? (1 point) 215 = 0.73n n = 0.73 × 215 0.73 = n × 215 215 = 73n

OpenStudy (anonymous):

2nd one you just change the % into a decimal and times the number to get that percentage of the number

OpenStudy (anonymous):

Thanks so much What percent of 94 is 23?

OpenStudy (anonymous):

Find the commission on a $1,250 sale with a commission rate of 5%

OpenStudy (anonymous):

23 = 94p do you get it?

OpenStudy (anonymous):

yes thanks :) Find the commission on a $1,250 sale with a commission rate of 5%

OpenStudy (anonymous):

n = 1'250 x 0.05